# Comprehending Knowledge Graphs with Large Language Models for Recommender Systems

### v2

@@ 5.1.1. Datasets
We conducted experiments on four real-world datasets, including three public datasets (MovieLens [^1] https://grouplens.org/datasets/movielens/, MIND [^2] https://msnews.github.io/, Last-FM [^3] https://grouplens.org/datasets/hetrec-2011/), and one industrial dataset (Fund). The statistics for these datasets are presented in Table 1. These datasets cover a wide range of application scenarios. Specifically, MovieLens is a well-established benchmark that collects movie ratings provided by users. MIND is a large-scale news recommendation dataset constructed from user click logs on Microsoft News. Last-FM is a well-known music recommendation dataset that includes user listening history and artist tags. The Fund dataset is sampled from the data of a large-scale online financial platform aiming to recommend funds for users. We adopt the similar setting as numerous previous studies (Xie et al., 2022; He et al., 2020), filtering out items and users with fewer than five interaction records. For each dataset, we randomly select 80% of each user’s historical interactions to form the training set, while the remaining 20% constitute the test set, following (He et al., 2020). From the training set, we further randomly select 10% of the interactions to create a validation set for tuning hyperparameters. Each observed user-item interaction is considered a positive instance, and we apply a negative sampling strategy by pairing it with one negative item that the user has not interacted with.

### v3

@@ 5.1.1. Datasets
We conducted experiments on four real-world datasets, including three public datasets (MovieLens [^1] https://grouplens.org/datasets/movielens/, MIND [^2] https://msnews.github.io/, Last-FM [^3] https://grouplens.org/datasets/hetrec-2011/), and one industrial dataset (Fund). The statistics for these datasets are presented in Table 1. These datasets cover a wide range of application scenarios. Specifically, MovieLens is a well-established benchmark that collects movie ratings provided by users. MIND is a large-scale news recommendation dataset constructed from user click logs on Microsoft News. Last-FM is a well-known music recommendation dataset that includes user listening history and artist tags. The Fund dataset is sampled from the data of a large-scale online financial platform aiming to recommend funds for users. We adopt the similar setting as numerous previous studies (Xie et al., 2022; He et al., 2020), filtering out items and users with fewer than five interaction records. For each dataset, we randomly select 80% of each user’s historical interactions to form the training set, while the remaining 20% constitute the test set, following (He et al., 2020). Each observed user-item interaction is considered a positive instance, and we apply a negative sampling strategy by pairing it with one negative item that the user has not interacted with.
@@ 6. Conclusion
7. Acknowledgments
This work was supported by the Early Career Scheme (No. CityU 21219323) and the General Research Fund (No. CityU 11220324) of the University Grants Committee (UGC), and the NSFC Young Scientists Fund (No. 9240127).